OpenStudy (anonymous):

What is the value of the function y = 3x - 1 when x = –1? A. –8 B. –4 C. 2 D. 6

pooja195 (pooja195):

solve it plug in -1

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

and ill stop harassing math

pooja195 (pooja195):

3(-1)-1 also dont threaten me its a study site answers come with work not for free

OpenStudy (anonymous):

no

pooja195 (pooja195):

3(-1)-1=?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

well done good job! :)

OpenStudy (anonymous):

What is the value of the function v = 3/2t - 5/2 when t = 3?

OpenStudy (anonymous):

As stated above, plug in the values.

pooja195 (pooja195):

Just plug in the value for x

OpenStudy (anonymous):

can u set it up

pooja195 (pooja195):

v = 3/2(3)- 5/2

OpenStudy (anonymous):

So its 2?

pooja195 (pooja195):

yes well done :)

OpenStudy (anonymous):

Find the value of the function when n = 6. p = 0.008n A. 0.048 B. 0.0042 C. 0.48 D. 6.008

OpenStudy (anonymous):

I do 0.008(6) right

pooja195 (pooja195):

yes :)

OpenStudy (anonymous):

hold on..

OpenStudy (anonymous):

thank u medal'd

OpenStudy (anonymous):

lates

pooja195 (pooja195):

your welcome :) and thanks